# Quantum synchronization between two strongly driven YIG spheres mediated via a microwave cavity

### v3

@@ IV Conclusion
Second, and more intriguingly, we observed the emergence of quantum synchronization between the coupled magnon modes (Fig. 4b). This phenomenon is quantitatively characterized by the synchronization measure $S_{q}^{\phi}$, which approaches unity under optimal conditions. In this regime, the individual quantum fluctuations of each mode become phase-locked, leading to coherent evolution despite the presence of decoherence and thermal noise. This indicates that the system dynamically suppresses relative quantum uncertainties, allowing for robust synchronization at the quantum level. In addition, we explored the effect of temperature on this delicate quantum behavior (Fig. 4b). We find that as thermal noise increases, the degree of quantum synchronization gradually decreases. This finding underscores the importance of low-temperature environments when trying to observe and maintain quantum synchronization in experiments.
@@ V ACKNOWLEDGMENTS
One of us (J.G.) acknowledges the financial support provided by the Department of Science and Technology-Innovation in Science Pursuit for Inspired Research (DST-INSPIRE), Government of India, during this work.
Appendix A Relevant coefficients in Eqs. (II.1)

### v4

@@ IV Conclusion
Second, and more intriguingly, we observed the emergence of quantum synchronization between the coupled magnon modes (Fig. 5). This phenomenon is quantitatively characterized by the synchronization measure $S_{q}^{\phi}$, which approaches unity under optimal conditions. In this regime, the individual quantum fluctuations of each mode become phase-locked, leading to coherent evolution despite the presence of decoherence and thermal noise. This indicates that the system dynamically suppresses relative quantum uncertainties, allowing for robust synchronization at the quantum level. In addition, we explored the effect of temperature on this delicate quantum behavior (Fig. 5). We find that as thermal noise increases, the degree of quantum synchronization gradually decreases. This finding underscores the importance of low-temperature environments when trying to observe and maintain quantum synchronization in experiments.
@@ V ACKNOWLEDGMENTS
One of us (J. Ghildiyal) acknowledges the financial support provided by the Department of Science and Technology-Innovation in Science Pursuit for Inspired Research (DST-INSPIRE), Government of India, through the fellowship DST/INSPIRE Fellowship/2019/IF190615 during this work.
Appendix A Relevant coefficients in Eqs. (II.2)